OpenStudy (anonymous):

C^2=a^2+b^2, solve for B

OpenStudy (anonymous):

subtract a^2 c^2-a^2=b^2 take the square root of each side \[\sqrt{c ^{2}-a ^{2}}\]=\[\sqrt{b ^{2}}\] b= \[\pm \sqrt{c ^{2}-a ^{2}}\]

OpenStudy (anonymous):

1 more, 1/r=1/r1+1/r2; r2 (solve for r2)

OpenStudy (anonymous):

is the r1 and r2 r^1 and r^2?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no nevermind

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

multiply each side by r2 r2/r=1/r1 multiply each side by r r2=r/r1

OpenStudy (anonymous):

thanks